On January 1, 2021, Nath-Langstrom Services, Inc., a computer software training firm, leased several computers under a two-year operating lease agreement from ComputerWorld Leasing, which routinely finances equipment for other firms at an annual interest rate of 4%. The contract calls for four rent payments of $19,500 each, payable semiannually on June 30 and December 31 each year. The computers were acquired by ComputerWorld at a cost of $109,000 and were expected to have a useful life of Five years with no residual value. Both firms record amortization and depreciation semiannually. (FV of $1, PV of $1, FVA of $1, PVA of $1, FVAD of $1 and PVAD of $1) (Use appropriate factor(s) from the tables.)

Required:
Prepare the appropriate entries for both the lessee and the lessor from the beginning of the lease through the end of 2021. (If no entry is required for a transaction/event, select "No journal entry required" in the first account field. (Round your intermediate calculations and final answers to the nearest whole dollar.)
Answer #1
In the books of Lessee (Nath-langstrom services Inc.)
Date Account titles and explanation Debit Credit
2021
Jan 1. Right of use of asset (Note:1) 70783
Lease payable 70783
(Lease liability recorded)
June 30. Interest expense (70783*4%*1/2) 1416
Lease payable (19500-1416) 18084
Cash 19500
(Lease payment made)
June 30. Amortization expense 18084
Right of use of asset 18084
(Amortization recorded)
Dec 31. Interest expense (70783-18084)*4%*1/2 1054
Lease payable (19500-1054) 18446
Cash 19500
(Lease payment made)
Dec 31. Amortization expense 18446
Right of use of asset 18446
(Amortization recorded)
Note:1
Lease rental=$ 19500
Discount factor=Interest rate=4%
Lease term=4 rental payments
Lease liability=Lease rental*Discount factor at 4% for 4 years=19500*3.629895=70782.95=$ 70783
In the books of Lessor (Computer world)
Date Account titles and explanation Debit Credit
Jan 1. No entry
(Since it is an operating lease)
June 30. Cash 19500
Lease revenue 19500
(Lease revenue recorded)
June 30. Depreciation expense (Note:2) 10900
Accumulated depreciation 10900
(Depreciation recorded)
Dec 31. Cash 19500
Lease revenue 19500
(Lease revenue recorded)
Dec 31. Depreciation expense (Note:2) 10900
Accumulated depreciation 10900
(Depreciation recorded)
Note:2
Depreciation expense for the year=(Cost-Residual value)/Useful life=(109000-0)/5=$ 21800
Depreciation expense for 6 months=21800/2=$ 10900